# A multi-objective mixed integer linear programming model for supply chain planning of 3D printing

### v2

@@ 1 Introduction
The increase in population and the growing need for natural resources such as water, energy, food, etc., to meet the needs of the rising population, have prompted the current generation to seek solutions for fulfilling the needs of future generations. Many of the current methods used to meet human needs are not sustainable and have adverse effects on the environment, economy, human health, etc. Therefore, many scientists and researchers are looking for ways to meet human needs that are renewable and have minimal long-term negative impacts on the environment and the life cycle. 3D printing, as one of the newest production systems in recent decades, has garnered much attention from researchers and industry professionals and is rapidly advancing. This new technology, utilizing minimal raw materials and energy, is considered the future of mass production systems. It can meet many human needs in industrial, medical, food, aerospace, and other fields sustainably and cost-effectively. The significance of this topic in these areas has led to extensive research for the commercialization and industrialization of this technology. Since production planning in supply chains is one of the significant issues for significantly reducing costs, production time, and so on, the production planning of 3D printers is crucial in the implementation of this technology in the industry. Many old-fashioned industrial companies can revamp their old production systems and benefit from the numerous advantages of 3D printing technology, such as high flexibility in product manufacturing, low cost, high accuracy, and so forth, by investing in this technology. 3D printers, with only digital information of the final products, can produce consumer goods in a short time and small quantities. This aspect can be beneficial for the companies in several ways. For example, products are manufactured exactly according to the specifications ordered by customers, which is a significant advantage in competitive markets. Additionally, the proximity of production centers using additive technology to final consumers reduces transportation costs. Optimal energy consumption, raw materials, and the need for fewer employees lower the final product price, which is a major competitive advantage. The significant question raised in this field is: How can we maximize the benefits of 3D printing technology? given the importance of production planning in manufacturing systems, how can we implement this effectively in these systems? It is important to note that 3D printers have only begun their industrialization process in recent years, as they still do not have high speeds for mass production and currently have fewer advantages compared to mass production systems. However, with the significant advancements in these machines over the past decade, it is expected that many organizations will replace conventional production systems with this system in the coming years. One of the best production approaches is for manufacturers to deliver customer orders exactly on time and essentially maintain zero inventory, which reduces the total cost of goods. The costs of purchasing, maintaining, and operating 3D printers are high because they have not yet reached the stage of mass production. Therefore, planning that optimizes the use of these devices and ensures timely delivery of orders will reduce production, maintenance, and lost profit costs. The main goal of this paper is to find such a plan for 3D printers. To address this problem, we proposed a multi-objective mixed integer linear programming (MOMILP) model. To the best of our knowledge, there exists no well-known work integrating the just-in-time concept of part manufacturing combined with part orientation optimization and maximization of 3D printer utilization. In particular, the orientation of parts to be produced by 3D printers can cause delays in order satisfaction and machine utilization rate. Our model includes the part orientation optimization and also maximization of 3D printer utilization. Indeed, the goal of our model is to maximize the utilization of machines and minimize the earliness or tardiness of parts manufacturing including part orientation optimization. Finally, we conducted several experiments on randomly generated data to validate our model. Several managerial insights have been concluded from the results, i.e., part orientation optimization can significantly reduce the earliness and tardiness of the production of the part given the available surface area of the printers. Similarity and discrepancies in dimensions and due dates of parts in a part are important in reducing the sum of earliness and tardiness of parts productions highlighting the significance of job allocation and sequencing.
@@ 2 Background
Rapid manufacturing: After a while, 3D printing technology began to be used for the direct production of final products. At this stage, the term rapid prototyping could no longer describe this new capability. The new term rapid manufacturing better described this capability.
@@ 2.1 History
3D printing technology began to grow rapidly in the late 1990s. However, this technology entered a competitive phase with traditional manufacturing after a two-decade delay and witnessed significant technical and entrepreneurial changes. Initially, this technology was used in industrial design markets before products reached mass production. This technology allowed manufacturers to produce parts in limited quantities much faster. With 3D printing, initial prototyping, which previously required several months or weeks with traditional manufacturing methods due to the need for molds and lengthy production lines, could be completed within hours or days. Moreover, 3D printers enabled designers to create parts with greater structural and geometric complexity. As a result of these advancements, improvements in quality, cost, and time have made 3D printing a leading method for prototyping [3].
@@ 2.2 The Fourth Industrial Revolution
[3, 6] reported that over two decades of 3D printer commercialization led to a cumulative annual growth rate of 25.4% in global revenue for 3D printing products in 2013. The growth rate rose by 27.4% from 2010 to 2012, achieving \$2.2 million in 2012. The sales of 3D printers grew by 19.3% in 2012, and the purchases of personal 3D printers saw a 46.3% rise in the same year, reflecting the growing popularity of 3D printers among both retailers and consumers. [7] conducted another study examining the factors for using 3D printers. They did a global survey with 330 participants, who were workers of organizations with over 100 staff, either utilizing 3D printers or intend to implement them. They showed that half of these organizations were using this technology in the early stages of production to improve their manufacturing processes. Interestingly, 6% of these individuals believed that the high cost of this technology was a barrier to its implementation, although companies currently using this technology believed that it had many benefits and that if the costs of implementation were lower and the technology improved, many companies would adopt this technology. Currently, over 67% of organizations are using this technology, and this percentage is expected to increase significantly in the future.
@@ 2.3.2 Medical Applications
Doctors at the Women and Children’s Hospital are using medical models produced by 3D printers to help children and patients with congenital heart diseases and complex conditions. Currently, 3D-printed medical models have become an integral part of treatment plans worldwide and are even used in hospitals for treating head and bone injuries. In recent years, hospitals relied on 2D scans for diagnosing and treating heart diseases, using MRI scans for heart diagnosis like most doctors. However, they now have basic structures for medical models. They can simply use data through 3D modeling software and 3D printers to construct the models within two hours. Researchers at MIT pioneered this process, and now doctors worldwide can have a medical model as part of their standard diagnostic package. They have proven the value of this method, which has been implemented in hospitals for women and children. Doctors have found that the benefits of medical 3D printing go beyond simple protection of the patient’s damaged heart. Doctors can cut the model and plan surgery based on the specific defects of each individual. This method is effectively applied in real surgeries, providing doctors with an opportunity to solve potential problems on 3D-printed medical models. This opportunity for surgical planning has led to advancements, meaning doctors spend less time in the operating room, patients recover faster, and significant cost savings are realized [10].
@@ 2.5.1 Production Planning
Production planning involves the strategic management of work and workloads within the production process, encompassing setting, controlling, and optimizing these activities. Factories employ both forward and backward planning techniques to allocate resources, plan human resources, schedule production tasks, and procure materials. Batch production planning specifically pertains to scheduling batch production processes. While scheduling is crucial for continuous traditional processes like refining [13], it is particularly critical for batch processes such as ”active pharmaceutical ingredients, biotechnology processes, and specialized chemical processes” [14]. Batch production planning shares methodologies with capacitated production planning, which addresses various production-related challenges [15]. Mathematical scheduling methods approach production planning as an optimization problem, aiming to minimize or maximize specific objectives under a set of constraints typically represented as equations. These objectives and constraints may involve binary variables, integers, and non-linear relationships. Constraint programming provides a systematic approach to solving such problems by treating them as a set of constraints, striving for efficient solutions in a timely manner [16]. For instance, [17] explored revenue management focusing on intricate pricing dynamics. They utilized a constrained MILP model to enhance revenue optimization, using product prices as decision variables while integrating capacity and consumer behavior within the model’s constraints.
@@ 2.5.3 Bin-packing problem
The bin-packing problem deals with placing parts of different sizes in a limited number of machines, using the minimum possible number of machines. Such problems are characterized by non-deterministic polynomial-time hardness [20]. This decision-making problem (deciding on parts to be placed in a specific number of machines) is one of the most difficult problems in this category and belongs to the NP-complete problem category [21]. Non-deterministic polynomial-time hardness in computational complexity theory is among the most difficult polynomial-time problems. A simple example of non-linear polynomial-time hardness problems is the subset sum problem [22]. Various types of these problems are discussed, such as linear packaging, weight-based packaging, cost-based packaging, and so on, which have applications in loading containers, trucks, and so on with weight constraints, etc.
@@ 2.5.4 Multi-objective Decision Making
$f_{i}(x^{1})\leq f_{i}(x^{2})\quad\forall i$
$f_{j}(x^{1})<f_{j}(x^{2})\quad\text{for at least one index}$
@@
$\min f_{j}(x)$
$s.t.$
$f_{i}(x)\leq\epsilon_{i}$ $\forall i=1,2,...,k\ \&\ j\neq i\text{ (}\epsilon_{i}=RL\text{)}$
@@ 3 Literature Review
Many steps are involved in transforming raw materials into final products within a production chain. Over recent decades, the emergence of 3D printing technology has revolutionized these steps in final product production, including design, planning, manufacturing, and sales. 3D printing enables unique products to reach customers in both small and large markets without the need for inventory or high production costs. Production planning is a framework that establishes production objectives, identifies necessary production resources, and formulates plans to achieve production goals on time. It anticipates future risks and prepares various scenarios to minimize waste. With the advent of 3D printing, production planning has transformed. One significant change is the reduction in the number of suppliers, as the amount of consumables in 3D printing is lower compared to traditional industrial production. In layer-by-layer production, parts produced are either final products or components of another assembly, ready for immediate use, whereas traditional production involves multiple stages to produce a single piece [28]. Quantitative research has focused on production planning and scheduling for 3D printing systems to optimize machine utilization. [29] proposed a method using MILP models to meet customer demand within delivery time windows, thereby reducing delays. Although their study does not address part orientation optimization, their primary objective is to minimize delivery delays. They also aim to minimize the time differential in delivering orders within a batch. [30] presented a mathematical model for scheduling SLM machines, an NP-hard problem aiming to minimize lateness and associated costs. To tackle the complexity, their paper proposes metaheuristic algorithms, specifically, Genetic Algorithms adapted for matrix handling. The capabilities of these algorithms are validated through a case study that includes a production plan with 30 parts. [31] tackled the challenge of printing uniquely configured parts across multiple machines with specific deadlines. The objective function was to minimize total time while adhering to all deadlines, utilizing a mixture of constrained bin packing, two-dimensional nesting, and job shop scheduling, framed through constraints and graph theory to create a formal model. [32] introduced a decision support model employing multi-objective optimization for managing batches of parts across several machines. The focus was on balancing operating costs, printer load, total tardiness, and the number of defective parts in the Fused Deposition Modeling (FDM) process. They validated the model by a case study involving automotive parts, and a trade-off analysis was conducted to examine conflicting objectives. [33] explored the maintenance and repair of 3D printers, noting that in the event of a sudden failure, the printer must restart the production process from the beginning after repairs. They proposed a stochastic preemptive-repeat scheduling model, establishing measures to minimize expected completion and flow times and solving an optimization problem to identify optimal job sizes for printing, specifically addressing dental 3D printing manufacturing challenges. [34] explored scheduling challenges for both single and multiple 3D printers by developing mathematical optimization models. They used MILP to assign parts to jobs on 3D printers, aiming to minimize makespan. Their approach considered various machine configurations such as single machines, parallel identical machines, and parallel non-identical machines. The models were solved using the CPLEX solver, providing optimal solutions within specific time limits (1800 and 2400 seconds). The study focused on scheduling issues involving parallel identical and non-identical Additive Manufacturing (AM) machines. [35] investigated the optimal allocation of parts to machines with diverse specifications, including production time, material cost per volume, processing time per unit height, setup time, maximum supported area, and height. They introduced a mathematical model implemented in CPLEX and devised heuristic methods (’best-fit’ and ’adapted best-fit’ rules) to achieve efficient solutions. The heuristics were explained step by step with examples, and numerical tests demonstrated their effectiveness in reducing 3D printing processing costs through efficient planning. [36] introduced a strategy for managing dynamic order acceptance and scheduling in Powder Bed Fusion (PBF) systems for on-demand production. They aimed to maximize average profit per unit time by integrating bin packing, batch processing, dynamic scheduling, and decision-making into a metaheuristic approach. Experimental studies validated the profitability of their method for complex, NP-hard scheduling problems. [37] developed methods to optimize the arrangement of multiple parts in 3D printing. They employed a ”3D printing feature-based orientation optimization” to improve production quality by adjusting part orientations. Their ”parallel nesting” algorithm used part projection profiles to maximize efficiency, reducing overall build time and costs. [38] addressed part grouping and scheduling challenges in 3D printing with non-identical parallel machines. They proposed grouping parts by material and optimizing build area usage within machine capacity constraints. Their approach aimed to minimize makespan using a MILP model and evaluated three metaheuristic algorithms for efficiency and scalability. [39] explored scheduling problems in metal 3D printing using multiple non-identical Selective Laser Melting (SLM) machines, which varied in dimensions, speed, and cost. The parts also differed in width, length, height, volume, release date, and due date. The aim was to minimize total tardiness while ensuring that parts were allocated to build platforms without overlapping. They proposed a genetic algorithm to efficiently address this complex scheduling issue. [40] tackled a novel challenge in 3D printing scheduling involving unrelated parallel machines. They considered practical constraints such as two-dimensional packing and varying part release times, compounded by multiple part orientation options that influence processing duration. Their approach utilized a MILP model to minimize makespan, recognizing the problem’s NP-hard complexity. For larger-scale instances, they recommended an Adaptive Large Neighborhood Search (ALNS) algorithm. [41] investigated the impact of build orientation policies on production time in 3D printing, focusing on mass customization models. They evaluated two primary strategies: the Laying Policy (LP), which reduces individual build times by minimizing part height but may increase overall completion times with more parts, and the Standing Policy (SP), which reduces job numbers by lowering base plane projections, although each job takes longer due to increased part height. Numerical experiments using Stereolithography (SLA) indicated that SP could effectively reduce completion time for approximately 40 parts. [42] addressed the scheduling of unrelated parallel batch processing machines in Additive Manufacturing (AM), integrating part orientation selection to minimize makespan. They proposed a MILP model and developed a simulated annealing algorithm with packing strategies, highlighting the benefits of considering multiple part orientations in machine scheduling compared to fixed orientation scenarios. [43] conducted a comprehensive analysis of nesting and scheduling challenges in 3D printing, using bibliometric and systematic review methods to map the field’s evolution. They focused on scheduling and allocating parts to 3D printers to minimize the earliness and tardiness of deliveries while maximizing machine utilization through optimal part orientations. Their novel Multi-Objective Mixed Integer Linear Programming (MOMILP) model was validated through extensive numerical experiments and sensitivity analyses. Based on this recent review paper and the reviewed literature above, the problem addressed in this paper is unique. We address scheduling and allocating parts to 3D printers, focusing on two objectives: minimizing earliness 10 and tardiness of part delivery and maximizing machine utilization by selecting the best part height, that is selecting the orientation of the part in the 3D printers. To this end, we propose a novel MOMILP model and analyze and validate the model through several numerical experiments and sensitivity analysis.

### v4

@@ 1 Introduction
The increase in population and the growing need for natural resources such as water, energy, food, etc., to meet the needs of the rising population, have prompted the current generation to seek solutions for fulfilling the needs of future generations.
Many of the current methods used to meet human needs are not sustainable and have adverse effects on the environment, economy, human health, etc. Therefore, many scientists and researchers are looking for ways to meet human needs that are renewable and have minimal long-term negative impacts on the environment and the life cycle.
3D printing, as one of the newest production systems in recent decades, has garnered much attention from researchers and industry professionals and is rapidly advancing. This new technology, utilizing minimal raw materials and energy, is considered the future of mass production systems. It can meet many human needs in industrial, medical, food, aerospace, and other fields sustainably and cost-effectively. The significance of this topic in these areas has led to extensive research for the commercialization and industrialization of this technology. Since production planning in supply chains is one of the significant issues for significantly reducing costs, production time, and so on, the production planning of 3D printers is crucial in the implementation of this technology in the industry.
Many old-fashioned industrial companies can revamp their old production systems and benefit from the numerous advantages of 3D printing technology, such as high flexibility in product manufacturing, low cost, high accuracy, and so forth, by investing in this technology. 3D printers, with only digital information of the final products, can produce consumer goods in a short time and small quantities. This aspect can be beneficial for the companies in several ways. For example, products are manufactured exactly according to the specifications ordered by customers, which is a significant advantage in competitive markets. Additionally, the proximity of production centers using additive technology to final consumers reduces transportation costs. Optimal energy consumption, raw materials, and the need for fewer employees lower the final product price, which is a major competitive advantage.
The significant question raised in this field is: How can we maximize the benefits of 3D printing technology? Given the importance of production planning in manufacturing systems, how can we implement this effectively in these systems? It is important to note that 3D printers have only begun their industrialization process in recent years, as they still do not have high speeds for mass production and currently have fewer advantages compared to mass production systems. However, with the significant advancements in these machines over the past decade, it is expected that many organizations will replace conventional production systems with this system in the coming years.
One of the best production approaches is for manufacturers to deliver customer orders exactly on time and essentially maintain zero inventory, which reduces the total cost of goods. The costs of purchasing, maintaining, and operating 3D printers are high because they have not yet reached the stage of mass production. Therefore, planning that optimizes the use of these devices and ensures the timely delivery of orders will reduce production, maintenance, and lost profit costs. The main goal of this paper is to find such a plan for 3D printers. To address this problem, we proposed a multi-objective mixed integer linear programming (MOMILP) model.
To the best of our knowledge, there exists no well-known work integrating the just-in-time concept of part manufacturing combined with part orientation optimization and maximization of 3D printer utilization. In particular, the orientation of parts to be produced by 3D printers can cause delays in order satisfaction and machine utilization rate. Our model includes the part orientation optimization and also the maximization of 3D printer utilization. Indeed, the goal of our model is to maximize the utilization of machines and minimize the earliness or tardiness of parts manufacturing, including part orientation optimization. Finally, we conducted several experiments on randomly generated data to validate our model. Several managerial insights have been concluded from the results, i.e., part orientation optimization can significantly reduce the earliness and tardiness of the production of the part given the available surface area of the printers. Similarities and discrepancies in dimensions and due dates of parts in a part are important in reducing the sum of earliness and tardiness of parts productions highlighting the significance of job allocation and sequencing.
@@ 2 Background
Rapid manufacturing: after a while, 3D printing technology began to be used for the direct production of final products. At this stage, the term rapid prototyping could no longer describe this new capability. The new term, rapid manufacturing, better describes this capability.
@@ 2.1 History
3D printing technology began to grow rapidly in the late 1990s. However, this technology entered a competitive phase with traditional manufacturing after a two-decade delay and witnessed significant technical and entrepreneurial changes.
Initially, this technology was used in industrial design markets before products reached mass production. This technology allowed manufacturers to produce parts in limited quantities much faster. With 3D printing, initial prototyping, which previously required several months or weeks with traditional manufacturing methods due to the need for molds and lengthy production lines, could be completed within hours or days. Moreover, 3D printers enabled designers to create parts with greater structural and geometric complexity. As a result of these advancements, improvements in quality, cost, and time have made 3D printing a leading method for prototyping [3].
@@ 2.2 The Fourth Industrial Revolution
[3, 6] reported that over two decades of 3D printer commercialization led to a cumulative annual growth rate of 25.4% in global revenue for 3D printing products in 2013. The growth rate rose by 27.4% from 2010 to 2012, achieving \$2.2 million in 2012. The sales of 3D printers grew by 19.3% in 2012, and the purchases of personal 3D printers saw a 46.3% rise in the same year, reflecting the growing popularity of 3D printers among both retailers and consumers.
[7] conducted another study examining the factors for using 3D printers. They did a global survey with 330 participants, who were workers of organizations with over 100 staff, either utilizing 3D printers or intend to implement them. They showed that half of these organizations were using this technology in the early stages of production to improve their manufacturing processes.
In particular, 6% of these individuals believed that the high cost of this technology was a barrier to its implementation, although companies currently using this technology believed that it had many benefits and that if the costs of implementation were lower and the technology improved, many companies would adopt this technology. Currently, over 67% of organizations are using this technology, and this percentage is expected to increase significantly in the future.
@@ 2.3.2 Medical Applications
Doctors at the Women and Children’s Hospital are using medical models produced by 3D printers to help children and patients with congenital heart diseases and complex conditions. Currently, 3D-printed medical models have become an integral part of treatment plans worldwide and are even used in hospitals for treating head and bone injuries. In recent years, hospitals have relied on 2D scans for diagnosing and treating heart diseases, using MRI scans for heart diagnosis, like most doctors. However, they now have basic structures for medical models. They can simply use data through 3D modeling software and 3D printers to construct the models within two hours. Researchers at MIT pioneered this process, and now doctors worldwide can have a medical model as part of their standard diagnostic package. They have proven the value of this method, which has been implemented in hospitals for women and children. Doctors have found that the benefits of medical 3D printing go beyond simple protection of the patient’s damaged heart. Doctors can cut the model and plan surgery based on the specific defects of each individual. This method is effectively applied in real surgeries, providing doctors with an opportunity to solve potential problems on 3D-printed medical models. This opportunity for surgical planning has led to advancements, meaning doctors spend less time in the operating room, patients recover faster, and significant cost savings are realized [10].
@@ 2.5.1 Production Planning
Production planning involves the strategic management of work and workloads within the production process, encompassing setting, controlling, and optimizing these activities. Factories employ both forward and backward planning techniques to allocate resources, plan human resources, schedule production tasks, and procure materials. Batch production planning specifically pertains to scheduling batch production processes.
While scheduling is crucial for continuous traditional processes like refining [13], it is particularly critical for batch processes such as ”active pharmaceutical ingredients, biotechnology processes, and specialized chemical processes” [14].
Batch production planning shares methodologies with capacitated production planning, which addresses various production-related challenges [15].
Mathematical scheduling methods approach production planning as an optimization problem, aiming to minimize or maximize specific objectives under a set of constraints typically represented as equations. These objectives and constraints may involve binary variables, integers, and non-linear relationships. Constraint programming provides a systematic approach to solving such problems by treating them as a set of constraints, striving for efficient solutions in a timely manner [16].
For instance, [17] explored revenue management focusing on intricate pricing dynamics. They utilized a constrained MILP model to enhance revenue optimization, using product prices as decision variables while integrating capacity and consumer behavior within the model’s constraints.
@@ 2.5.3 Bin-packing problem
The bin-packing problem deals with placing parts of different sizes in a limited number of machines, using the minimum possible number of machines. Such problems are characterized by non-deterministic polynomial-time hardness [20]. This decision-making problem (deciding on parts to be placed in a specific number of machines) is one of the most difficult problems in this category and belongs to the NP-complete problem category [21].
Non-deterministic polynomial-time hardness in computational complexity theory is among the most difficult polynomial-time problems. A simple example of non-linear polynomial-time hardness problems is the subset sum problem [22].
Various types of these problems are discussed, such as linear packaging, weight-based packaging, cost-based packaging, and so on, which have applications in loading containers, trucks, and so on with weight constraints, etc.
@@ 2.5.4 Multi-objective Decision Making
$f_{i}(x^{1})\leq f_{i}(x^{2}),\quad\forall i$
$f_{j}(x^{1})<f_{j}(x^{2}),\quad\text{for at least one index}$
@@
$\min_{x}$ $\ f_{j}(x)$
$s.t.$ $\ f_{i}(x)\leq\epsilon_{i}$ $\forall i,j=1,2,...,k\ \&\ j\neq i\text{ (}\epsilon_{i}=RL\text{)}$
@@ 3 Literature Review
Many steps are involved in transforming raw materials into final products within a production chain. Over recent decades, the emergence of 3D printing technology has revolutionized these steps in final product production, including design, planning, manufacturing, and sales. 3D printing enables the creation of unique products that can reach customers in both small and large markets, eliminating the need for inventory and reducing high production costs.
Production planning is a framework that establishes production objectives, identifies necessary production resources, and formulates plans to achieve production goals on time. It anticipates future risks and prepares various scenarios to minimize waste. With the advent of 3D printing, production planning has undergone a significant transformation. One significant change is the reduction in the number of suppliers, as the amount of consumables in 3D printing is lower compared to traditional industrial production.
In layer-by-layer production, parts produced are either final products or components of another assembly, ready for immediate use, whereas traditional production involves multiple stages to produce a single piece [28]. Quantitative research has focused on production planning and scheduling for 3D printing systems to optimize machine utilization.
[29] proposed a method using MILP models to meet customer demand within delivery time windows, thereby reducing delays. Although their study does not address part orientation optimization, their primary objective is to minimize delivery delays. They also aim to minimize the time differential in delivering orders within a batch.
[30] presented a mathematical model for scheduling SLM machines, an NP-hard problem aiming to minimize lateness and associated costs. To tackle the complexity, their paper proposes metaheuristic algorithms, specifically, Genetic Algorithms adapted for matrix handling. The capabilities of these algorithms are validated through a case study that includes a production plan with 30 parts.
[31] tackled the challenge of printing uniquely configured parts across multiple machines with specific deadlines. The objective function was to minimize total time while adhering to all deadlines, utilizing a mixture of constrained bin packing, two-dimensional nesting, and job shop scheduling, framed through constraints and graph theory to create a formal model.
[32] introduced a decision support model employing multi-objective optimization for managing batches of parts across several machines. The focus was on balancing operating costs, printer load, total tardiness, and the number of defective parts in the Fused Deposition Modeling (FDM) process. They validated the model by a case study involving automotive parts, and a trade-off analysis was conducted to examine conflicting objectives.
[33] explored the maintenance and repair of 3D printers, noting that in the event of a sudden failure, the printer must restart the production process from the beginning after repairs. They proposed a stochastic preemptive-repeat scheduling model, establishing measures to minimize expected completion and flow times and solving an optimization problem to identify optimal job sizes for printing, specifically addressing dental 3D printing manufacturing challenges.
[34] explored scheduling challenges for both single and multiple 3D printers by developing mathematical optimization models. They used MILP to assign parts to jobs on 3D printers, aiming to minimize makespan. Their approach considered various machine configurations such as single machines, parallel identical machines, and parallel non-identical machines. The models were solved using the CPLEX solver, providing optimal solutions within specific time limits (1800 and 2400 seconds). The study focused on scheduling issues involving parallel identical and non-identical Additive Manufacturing (AM) machines.
[35] investigated the optimal allocation of parts to machines with diverse specifications, including production time, material cost per volume, processing time per unit height, setup time, maximum supported area, and height. They introduced a mathematical model implemented in CPLEX and devised heuristic methods (’best-fit’ and ’adapted best-fit’ rules) to achieve efficient solutions. The heuristics were explained step by step with examples, and numerical tests demonstrated their effectiveness in reducing 3D printing processing costs through efficient planning.
[36] introduced a strategy for managing dynamic order acceptance and scheduling in Powder Bed Fusion systems for on-demand production. They aimed to maximize average profit per unit time by integrating bin packing, batch processing, dynamic scheduling, and decision-making into a metaheuristic approach. Experimental studies validated the profitability of their method for complex, NP-hard scheduling problems.
[37] developed methods to optimize the arrangement of multiple parts in 3D printing. They employed a ”3D printing feature-based orientation optimization” to improve production quality by adjusting part orientations. Their ”parallel nesting” algorithm used part projection profiles to maximize efficiency, reducing overall build time and costs.
[38] addressed part grouping and scheduling challenges in 3D printing with non-identical parallel machines. They proposed grouping parts by material and optimizing build area usage within machine capacity constraints. Their approach aimed to minimize makespan using an MILP model and evaluated three metaheuristic algorithms for efficiency and scalability.
[39] explored scheduling problems in metal 3D printing using multiple non-identical Selective Laser Melting (SLM) machines, which varied in dimensions, speed, and cost. The parts also differed in width, length, height, volume, release date, and due date. The aim was to minimize total tardiness while ensuring that parts were allocated to build platforms without overlapping. They proposed a genetic algorithm to efficiently address this complex scheduling issue.
[40] tackled a novel challenge in 3D printing scheduling involving unrelated parallel machines. They considered practical constraints such as two-dimensional packing and varying part release times, compounded by multiple part orientation options that influence processing duration. Their approach utilized an MILP model to minimize makespan, recognizing the problem’s NP-hard complexity. For larger-scale instances, they recommended an Adaptive Large Neighborhood Search algorithm.
[41] investigated the impact of build orientation policies on production time in 3D printing, focusing on mass customization models. They evaluated two primary strategies: the Laying Policy, which reduces individual build times by minimizing part height but may increase overall completion times with more parts, and the Standing Policy (SP), which reduces job numbers by lowering base plane projections, although each job takes longer due to increased part height. Numerical experiments using Stereolithography indicated that SP could effectively reduce completion time for approximately 40 parts.
[42] addressed the scheduling of unrelated parallel batch processing machines in AM, integrating part orientation selection to minimize makespan. They proposed a MILP model and developed a simulated annealing algorithm with packing strategies, highlighting the benefits of considering multiple part orientations in machine scheduling compared to fixed orientation scenarios.
[43] conducted a comprehensive analysis of nesting and scheduling challenges in 3D printing, using bibliometric and systematic review methods to map the field’s evolution. They focused on scheduling and allocating parts to 3D printers to minimize the earliness and tardiness of deliveries while maximizing machine utilization through optimal part orientations. Their novel MOMILP model was validated through extensive numerical experiments and sensitivity analyses. Based on this recent review paper and the reviewed literature above, the problem addressed in this paper is unique.
We address scheduling and allocating parts to 3D printers, focusing on two objectives: minimizing earliness 10 and the tardiness of part delivery, and maximizing machine utilization by selecting the best part height, that is, selecting the orientation of the part in the 3D printers. To this end, we propose a novel MOMILP model and analyze and validate the model through several numerical experiments and sensitivity analysis.